# Anomaly-induced vanishing of brane partition functions

### v1

@@ 2.5 Insertion of sources
evaluates to zero Witten:1999vg . This is no longer true when $\mathcal{O}=\mathcal{O}[\check{A}]$ is an operator which also depends on the background gauge field $\check{A}\in\check{Z}^{p+2}(X^{d})$, in a way such that its insertion enforces the modified basepoint anomaly
@@
It is important to note that the basepoint anomaly $\mathcal{A}[\check{A},0,\check{a}]$ is an “anomaly” not in the sense that the resultant theory is inconsistent, but simply that the corresponding partition function vanishes Hsieh:2020jpj ; Yonekura:2024bvh . As a result, it would not contribute to the path integral if one were to promote $\check{A}$ to a dynamical gauge field, e.g. in a gravitational theory where all symmetries are expected to be gauged Harlow:2018tng . Particularly, our formalism provides a unified approach to a variety of examples in string/M/F-theory, some of which were previously known, where certain “tadpole constraints” or “consistency conditions” are required to hold.
@@ 3.1 Generalised Maxwell theory
These two global symmetries have a mixed anomaly given by the product, $(-1)^{d-p}\underline{\check{C}}\star\underline{\check{B}}$, of differential cocycles Hsieh:2020jpj , or more specifically,
@@ 3.1.1 Derivation with the mapping torus
i.e. the gauge fields $\check{B},\check{C}$ are topologically trivial, hence proving the claim in Hsieh:2020jpj .
@@ 3.1.3 Ordering (un)ambiguity
As was noted in Hopkins:2002rd ; Hsieh:2020jpj ; GarciaEtxebarria:2024fuk , unlike differential cohomology classes which are graded-commutative, the following two products of differential cocycles,
@@ 3.1.4 Wilson and ’t Hooft operators
We discussed in Section 1 that the insertion of a Wilson or ’t Hooft operator (i.e. the charged operators of Maxwell theory) generally leads to a vanishing correlator, but it becomes non-zero precisely when the conditions above are satisfied, which reflects the fact that Maxwell theory equipped with non-trivial cohomology classes $\mathsf{H}_{p+2}]_{\mathbb{Z}},[\mathsf{G}_{d-p}]_{\mathbb{Z}}$ is a perfectly well-defined theory. Despite having a vanishing partition function, one should not interpret such a theory as being inconsistent, as was pointed out by Yonekura:2024bvh .
@@ 3.2 Freed-Witten anomaly
where $w_{2}\in C^{2}(X^{4};\mathbb{R})$ is an uplift of the $\mathbb{Z}_{2}$-valued second Stiefel-Whitney class of $X^{4}$ to a real cochain Hsieh:2020jpj , via the following diagram.
@@
As argued by Kapustin:2014tfa ; Thorngren:2014pza ; Wang:2018qoy ; Hsieh:2020jpj ; Yonekura:2024bvh , the anomaly theory of 4d Maxwell theory with a $\text{Spin}^{c}$ gauge field is obtained effectively by replacing $\underline{\check{B}}\to\underline{\check{B}}+\underline{\check{w}}$ and $\underline{\check{C}}\to\underline{\check{C}}+\underline{\check{w}}$ in (58). More explicitly, we have
@@
Moreover, since $\mathsf{W}_{3}]_{\mathbb{Z}}=\beta([w_{2}]_{\mathbb{R}/\mathbb{Z}})$ with $w_{2}]_{\mathbb{R}/\mathbb{Z}}$ being 2-torsion, by exactness it must be that $2[\mathsf{H}_{3}]_{\mathbb{Z}}=2[\mathsf{G}_{3}]_{\mathbb{Z}}=0$ as well. [^18] See a similar consistency condition recently derived by Kim:2025fpz on 5d $\mathcal{N}=1$ $\text{SU}(2)$ super Yang-Mills theory on $X^{4}\times S^{1}$, wherein its relation to K-theoretic Donaldson invariants was discussed.
This recovers the celebrated Freed-Witten anomaly cancellation condition for D-branes in Type IIB string theory Freed:1999vc , as well as its S-dual counterpart Diaconescu:2000wy ; Yonekura:2024bvh . Note that all orientable 4-manifolds admit $\text{Spin}^{c}$ structures Teichner:1994abc , in which case $\mathsf{W}_{3}]_{\mathbb{Z}}=0$. Again we stress that a violation of (82) does not imply that placing the D-brane on such a background is prohibited, but rather that its partition function is necessarily vanishing.
@@
Consequently, even if the manifold is $\text{spin}^{c}$, one may still be able to have a non-vanishing D3-brane partition function in a background with non-trivial $\mathsf{H}_{3}]_{\mathbb{Z}}$ or $\mathsf{G}_{3}]_{\mathbb{Z}}$, as long as the appropriate sources are inserted Witten:1998xy ; Maldacena_2001 .
@@ 3.3 Dijkgraaf-Witten theory
With some minor modifications, we can apply our earlier results to theories whose anomaly is given by a Dijkgraaf-Witten theory Dijkgraaf:1989pz with $\mathbb{Z}_{N}$ gauge fields. One way to model such discrete gauge fields with differential cocycles is to uplift them to real cochains, analogously to what we did in the diagram (79). Alternatively, we can directly define such a gauge field as a doublet, i.e.
@@
Let us review the construction of a 4d $\mathfrak{u}(N)$ gauge theory, making use of the precriptions by Kapustin:2014gua ; Gaiotto:2014kfa ; Gaiotto:2017yup . We start with a 4d Yang-Mills theory where the global form of its gauge group is $G=\text{SU}(N)\times_{\mathbb{Z}_{k}}\mathrm{U}(1)$ for some generic divisor $k$ of $N$, whose connection can be parametrised as
@@
It is important to note that this is not a statement for the $\text{SU}(N)/\mathbb{Z}_{k}$ gauge theory. The $\mathrm{U}(1)$ factor in the gauge group actually matters. In fact, the cohomology classes $B_{2}]_{\mathbb{Z}_{k}}\in H^{2}(X^{4};\mathbb{Z}_{k})$ and $\mathsf{H}_{3}]_{\mathbb{Z}}\in H^{3}(X^{4};\mathbb{Z})$ play rather different roles depending on whether we are working with an $\text{SU}(N)/\mathbb{Z}_{k}$ gauge theory or an $\text{SU}(N)\times_{\mathbb{Z}_{k}}\mathrm{U}(1)$ gauge theory. Particularly, in the former case, a non-trivial $B_{2}]_{\mathbb{Z}_{k}}$ is an obstruction to lift to an $\text{SU}(N)$ gauge theory, while a non-trivial $\mathsf{H}_{3}]_{\mathbb{Z}}$ is an obstruction to lift to an $\text{SU}(N)\times_{\mathbb{Z}_{k}}\mathrm{U}(1)$ gauge theory. [^20] For this reason, $B_{2}]_{\mathbb{Z}_{k}}$ is sometimes referred to in the physics literature as the second Stiefel-Whitney class of the $\text{SU}(N)/\mathbb{Z}_{k}$ gauge bundle Gaiotto:2017yup ; Hsin:2018vcg ; Brennan:2023vsa . See Table 1 for a summary.
@@ 3.5 Coincident D-branes
It was observed by Kapustin in Kapustin:1999di that the Freed-Witten anomaly cancellation condition (82) receives a correction when dealing with a stack of coincident D-branes. In the case of D3-branes, this result can indeed be recast in our language as follows.
@@
When we take $k=N$, the first relation in (105) matches with the original result in Kapustin:1999di . By S-duality, we expect the second relation to hold as well. Nevertheless, our interpretations of these conditions are slightly different from Kapustin:1999di . The concern therein was whether the Chan-Paton structure group can be lifted from $\text{SU}(N)/\mathbb{Z}_{N}$ to $\mathrm{U}(N)$, but here we interpret the result as whether the partition function of the coincident D3-branes is vanishing or not. For example, suppose $\mathsf{W}_{3}]_{\mathbb{Z}}=0$, then the partition function is non-vanishing only if the characteristic class $\beta_{k}([\zeta_{2}^{B}])$ of the Chan-Paton bundle cancels out (the pullback of) $\mathsf{H}_{3}]_{\mathbb{Z}}$, and likewise for $\beta_{k}([\zeta_{2}^{C}])$ and $\mathsf{G}_{3}]_{\mathbb{Z}}$.
@@ 4 Other finite symmetry examples
It was shown in Hsin:2018vcg that any 3d topological quantum field theory (TQFT) with a 1-form $\mathbb{Z}_{N}$ global symmetry admits a factorisation, assuming $\gcd(N,p)=1$ for simplicity,
@@
where $\mathcal{A}^{N,p}$ is called a minimal Abelian TQFT, and $\mathcal{T}^{\prime}$ is a decoupled sector. In particular, the former consists of line operators charged under the $\mathbb{Z}_{N}$ symmetry, and have non-trivial braiding statistics labeled by an integer $p$, which takes values in $\mathbb{Z}_{N}$ on spin manifolds. [^21] For manifolds which are not spin (but orientable), $p\in\mathbb{Z}_{2N-1}$ for $N$ even and $p\in\mathbb{Z}_{2N-2}$ for $N$ odd. To be concrete, we focus on spin manifolds in this discussion. As an aside, the minimal TQFT plays a crucial role in the construction of non-invertible symmetry defects in 4d QFTs (see, e.g. Kaidi:2021xfk ; Choi:2022zal ).
@@
where $\widehat{\underline{B}}_{2}\in H^{2}(Y^{4};\mathbb{Z})$. Equivalently, as was pointed out by Karasik:2022kkq , the minimal TQFT $\mathcal{A}^{N,p}$ has a non-vanishing partition function only if the flux of the background connection through every 2-cycle of $X^{3}$ is a multiple of NN. [^22] A similar observation was made in Appendix A of Choi:2022zal .
@@
where the cohomology operation $\mathfrak{B}_{p}:H^{i}(-;\mathbb{Z}_{pm})\to H^{pi}(-;\mathbb{Z}_{p^{2}m})$ for any prime $p$ and any $m\in\mathbb{Z}^{+}$ is the Pontryagin $p$-th power b8effd3d-9c0a-396c-a024-ee4cd0bee32e . The case of $p=2$ and $m=1$ corresponds to the standard Pontryagin square operation Whitehead1949 . One of its axioms is that
@@
where $\cup_{i}:C^{|x|}(-;\mathbb{Z}_{pm})\times C^{|y|}(-;\mathbb{Z}_{pm})\to C^{|x|+|y|-i}(-;\mathbb{Z}_{p^{2}m})$ denotes the cup-$i$ product 4fc46540-015c-3832-bc1c-ffd13cef2752 , satisfying
@@ 4.2 Vanishing of the RR partition function in 2d
Consider a free Dirac fermion in $2$ dimensions. Like the $4d$ system, this has vector and axial $\mathrm{U}(1)$ symmetries with a mixed anomaly, leading to a vanishing condition similar to the one in the case of the ABJ anomaly discussed in the introduction. However, in this case, there is an additional structure that we can exploit, as discussed in Karch_2019 .
To define a fermion theory on a Riemann surface $\Sigma^{2}$, we need to specify a Spin structure. One can think of this as specifying periodic (R) or antiperiodic (NS) boundary conditions to each cycle on $\Sigma^{2}$. A classical result, derived e.g. in Hori2003MirrorSymmetry , is that the partition function of a free fermion on the torus with RR Spin structure vanishes. When quantising the theory, one finds that there is a fermionic zero mode, which causes the path integral to vanish. On general $\Sigma^{2}$, the appearance of such zero modes, and therefore the vanishing, is controlled by the mod $2$ index, i.e. the number of zero modes modulo $2$. In $2d$, this index can be non-zero and is a topological invariant. A theorem due to Atiyah Atiyah1971 states that
@@
where $\mathcal{I}(\rho_{\Sigma^{2}})$ is the mod $2$ index associated to the Spin structure $\rho_{\Sigma^{2}}$. To understand the right-hand-side, recall that the first homology group with $\mathbb{Z}_{2}$ coefficients of the genus $g$ Riemann surface is given by
@@
We may define the intersection form for two cycles represented by closed curves $\gamma$, $\delta$ in $\Sigma^{2}$: [^24] We are making the additional technical assumption that $\gamma$ and $\delta$ only intersect transversely. Intuitively, we can always deform them slightly to satisfy this assumption without changing the homology classes they represent. A proof can be found in standard references for differential topology, such as guillemin2010differential .
@@
for any $a,b\in H_{1}(\Sigma^{2};\mathbb{Z}_{2})$. It can be shown that there is a bijection between quadratic refinements of the intersection form and spin structures on $\Sigma^{2}$ johnson1980spin . Once we have such a quadratic refinement of the intersection form, we can define its associated $\mathbb{Z}_{2}$-valued Arf invariant,
@@
where $\gamma^{3}$ is the chirality matrixKarch_2019 ; Delmastro:2021xox . By general arguments witten2020anomalyinflowetainvariant , the anomaly theory of this system is given by the $\eta$ invariant of a real Dirac operator in $3d$ coupled to this $\mathbb{Z}_{2}$-background. Potential anomalies of this symmetry are classified by $\Omega^{\text{Spin}}_{3}(B\mathbb{Z}_{2})=\mathbb{Z}_{8}$. This group is generated by $\mathbb{RP}^{3}$ with non-trivial $\mathbb{Z}_{2}$ bundle. As pointed out in Tachikawa:2018njr , the relevant $\eta$-invariant evaluates to $\pm 1/8$ on this generator, so our $(-1)^{F_{L}}$-symmetry is indeed anomalous. To see the connection between this $\eta$-invariant and the Arf invariant discussed above we need to take one additional mathematical detour.
A quadratic enhancement of the intersection form is a function $\tilde{q}:H_{1}(\Sigma^{2};\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{4}$ such that $\tilde{q}(a+b)=\tilde{q}(a)+\tilde{q}(b)+2a\cdot b$. A Pin- structure gives rise to such a quadratic enhancement. Since every closed $2$-manifold admits a Pin- structure Kirby_Taylor_1991 but only orientable surfaces admit a Spin structure, this provides a generalisation of the quadratic refinement above. On spin surfaces, the quadratic enhancement $\tilde{q}$ is related to the quadratic refinement $q$ as
@@
We omit further details of this construction and refer the interested reader to Kirby_Taylor_1991 .
@@
This theory has been discussed in the physics literature, see e.g. Guo:2018vij ; Grigoletto:2021zyv . The Poincaré dual of $x$ can be represented by a closed surface embedded in $\mathcal{M}^{3}$. Since every closed surface is pin-, the ambient Spin structure $\rho$ induces a Pin- structure $s$ on PD(x)(x). Consider our generator of $\Omega^{\text{Spin}}_{3}(B\mathbb{Z}_{2})$. The Poincaré dual of the non-trivial element $x\in H^{1}(\mathbb{RP}^{3};\mathbb{Z}_{2})$ is an embedded $\mathbb{RP}^{2}$. Depending on the choice of Pin- structure, $\beta(\mathbb{RP}^{2},s)=\pm 1\in\mathbb{Z}_{8}$. Therefore, the partition function of the theory $\mathcal{A}$ agrees with the exponentiated $\eta$-invariant on the generator of $\Omega^{\text{Spin}}_{3}(B\mathbb{Z}_{2})$. Since both quantities are bordism invariants, we deduce that they agree on all closed spin $3$-manifolds with $\mathbb{Z}_{2}$ bundle. That is, $\mathcal{A}$ is a presentation of the anomaly theory of our Majorana fermion.

### v3

@@ 2.5 Insertion of sources
evaluates to zero Witten (2000). This is no longer true when $\mathcal{O}=\mathcal{O}[\check{A}]$ is an operator which also depends on the background gauge field $\check{A}\in\check{Z}^{p+2}(X^{d})$, in a way such that its insertion enforces the modified basepoint anomaly
@@
It is important to note that the basepoint anomaly $\mathcal{A}[\check{A},0,\check{a}]$ is an “anomaly” not in the sense that the resultant theory is inconsistent, but simply that the corresponding partition function vanishes Hsieh et al. (2022); Yonekura (2024). As a result, it would not contribute to the path integral if one were to promote $\check{A}$ to a dynamical gauge field, e.g. in a gravitational theory where all symmetries are expected to be gauged Harlow and Ooguri (2021). Particularly, our formalism provides a unified approach to a variety of examples in string/M/F-theory, some of which were previously known, where certain “tadpole constraints” or “consistency conditions” are required to hold.
@@ 3.1 Generalised Maxwell theory
These two global symmetries have a mixed anomaly given by the product, $(-1)^{d-p}\underline{\check{C}}\star\underline{\check{B}}$, of differential cocycles Hsieh et al. (2022), or more specifically,
@@ 3.1.1 Derivation with the mapping torus
i.e. the gauge fields $\check{B},\check{C}$ are topologically trivial, hence proving the claim in Hsieh et al. (2022).
@@ 3.1.3 Ordering (un)ambiguity
As was noted in Hopkins and Singer (2005); Hsieh et al. (2022); García Etxebarria and Hosseini (2025), unlike differential cohomology classes which are graded-commutative, the following two products of differential cocycles,
@@ 3.1.4 Wilson and ’t Hooft operators
We discussed in Section 1 that the insertion of a Wilson or ’t Hooft operator (i.e. the charged operators of Maxwell theory) generally leads to a vanishing correlator, but it becomes non-zero precisely when the conditions above are satisfied, which reflects the fact that Maxwell theory equipped with non-trivial cohomology classes $\mathsf{H}_{p+2}]_{\mathbb{Z}},[\mathsf{G}_{d-p}]_{\mathbb{Z}}$ is a perfectly well-defined theory. Despite having a vanishing partition function, one should not interpret such a theory as being inconsistent, as was pointed out by Yonekura (2024).
@@ 3.2 Freed-Witten anomaly
where $w_{2}\in C^{2}(X^{4};\mathbb{R})$ is an uplift of the $\mathbb{Z}_{2}$-valued second Stiefel-Whitney class of $X^{4}$ to a real cochain Hsieh et al. (2022), via the following diagram.
@@
As argued by Kapustin (2014); Thorngren (2015); Wang et al. (2019); Hsieh et al. (2022); Yonekura (2024), the anomaly theory of 4d Maxwell theory with a $\text{Spin}^{c}$ gauge field is obtained effectively by replacing $\underline{\check{B}}\to\underline{\check{B}}+\underline{\check{w}}$ and $\underline{\check{C}}\to\underline{\check{C}}+\underline{\check{w}}$ in (58). More explicitly, we have
@@
Moreover, since $\mathsf{W}_{3}]_{\mathbb{Z}}=\beta([w_{2}]_{\mathbb{R}/\mathbb{Z}})$ with $w_{2}]_{\mathbb{R}/\mathbb{Z}}$ being 2-torsion, by exactness it must be that $2[\mathsf{H}_{3}]_{\mathbb{Z}}=2[\mathsf{G}_{3}]_{\mathbb{Z}}=0$ as well. [^18] See a similar consistency condition recently derived by Kim et al. (2025) on 5d $\mathcal{N}=1$ $\text{SU}(2)$ super Yang-Mills theory on $X^{4}\times S^{1}$, wherein its relation to K-theoretic Donaldson invariants was discussed.
This recovers the celebrated Freed-Witten anomaly cancellation condition for D-branes in Type IIB string theory Freed and Witten (1999), as well as its S-dual counterpart Diaconescu et al. (2003a); Yonekura (2024). Note that all orientable 4-manifolds admit $\text{Spin}^{c}$ structures Teichner and Vogt (1994), in which case $\mathsf{W}_{3}]_{\mathbb{Z}}=0$. Again we stress that a violation of (82) does not imply that placing the D-brane on such a background is prohibited, but rather that its partition function is necessarily vanishing.
@@
Consequently, even if the manifold is $\text{spin}^{c}$, one may still be able to have a non-vanishing D3-brane partition function in a background with non-trivial $\mathsf{H}_{3}]_{\mathbb{Z}}$ or $\mathsf{G}_{3}]_{\mathbb{Z}}$, as long as the appropriate sources are inserted Witten (1998); Maldacena and Seiberg (2001).
@@ 3.3 Dijkgraaf-Witten theory
With some minor modifications, we can apply our earlier results to theories whose anomaly is given by a Dijkgraaf-Witten theory Dijkgraaf and Witten (1990) with $\mathbb{Z}_{N}$ gauge fields. One way to model such discrete gauge fields with differential cocycles is to uplift them to real cochains, analogously to what we did in the diagram (79). Alternatively, we can directly define such a gauge field as a doublet, i.e.
@@
Let us review the construction of a 4d $\mathfrak{u}(N)$ gauge theory, making use of the precriptions by Kapustin and Seiberg (2014); Gaiotto et al. (2015, 2017). We start with a 4d Yang-Mills theory where the global form of its gauge group is $G=\text{SU}(N)\times_{\mathbb{Z}_{k}}\mathrm{U}(1)$ for some generic divisor $k$ of $N$, whose connection can be parametrised as
@@
It is important to note that this is not a statement for the $\text{SU}(N)/\mathbb{Z}_{k}$ gauge theory. The $\mathrm{U}(1)$ factor in the gauge group actually matters. In fact, the cohomology classes $B_{2}]_{\mathbb{Z}_{k}}\in H^{2}(X^{4};\mathbb{Z}_{k})$ and $\mathsf{H}_{3}]_{\mathbb{Z}}\in H^{3}(X^{4};\mathbb{Z})$ play rather different roles depending on whether we are working with an $\text{SU}(N)/\mathbb{Z}_{k}$ gauge theory or an $\text{SU}(N)\times_{\mathbb{Z}_{k}}\mathrm{U}(1)$ gauge theory. Particularly, in the former case, a non-trivial $B_{2}]_{\mathbb{Z}_{k}}$ is an obstruction to lift to an $\text{SU}(N)$ gauge theory, while a non-trivial $\mathsf{H}_{3}]_{\mathbb{Z}}$ is an obstruction to lift to an $\text{SU}(N)\times_{\mathbb{Z}_{k}}\mathrm{U}(1)$ gauge theory. [^20] For this reason, $B_{2}]_{\mathbb{Z}_{k}}$ is sometimes referred to in the physics literature as the second Stiefel-Whitney class of the $\text{SU}(N)/\mathbb{Z}_{k}$ gauge bundle Gaiotto et al. (2017); Hsin et al. (2019); Brennan and Intriligator (2024). See Table 1 for a summary.
@@ 3.5 Coincident D-branes
It was observed by Kapustin in Kapustin (2000) that the Freed-Witten anomaly cancellation condition (82) receives a correction when dealing with a stack of coincident D-branes. In the case of D3-branes, this result can indeed be recast in our language as follows.
@@
When we take $k=N$, the first relation in (105) matches with the original result in Kapustin (2000). By S-duality, we expect the second relation to hold as well. Nevertheless, our interpretations of these conditions are slightly different from Kapustin (2000). The concern therein was whether the Chan-Paton structure group can be lifted from $\text{SU}(N)/\mathbb{Z}_{N}$ to $\mathrm{U}(N)$, but here we interpret the result as whether the partition function of the coincident D3-branes is vanishing or not. For example, suppose $\mathsf{W}_{3}]_{\mathbb{Z}}=0$, then the partition function is non-vanishing only if the characteristic class $\beta_{k}([\zeta_{2}^{B}])$ of the Chan-Paton bundle cancels out (the pullback of) $\mathsf{H}_{3}]_{\mathbb{Z}}$, and likewise for $\beta_{k}([\zeta_{2}^{C}])$ and $\mathsf{G}_{3}]_{\mathbb{Z}}$.
@@ 4 Other finite symmetry examples
It was shown in Hsin et al. (2019) that any 3d topological quantum field theory (TQFT) with a 1-form $\mathbb{Z}_{N}$ global symmetry admits a factorisation, assuming $\gcd(N,p)=1$ for simplicity,
@@
where $\mathcal{A}^{N,p}$ is called a minimal Abelian TQFT, and $\mathcal{T}^{\prime}$ is a decoupled sector. In particular, the former consists of line operators charged under the $\mathbb{Z}_{N}$ symmetry, and have non-trivial braiding statistics labeled by an integer $p$, which takes values in $\mathbb{Z}_{N}$ on spin manifolds. [^21] For manifolds which are not spin (but orientable), $p\in\mathbb{Z}_{2N-1}$ for $N$ even and $p\in\mathbb{Z}_{2N-2}$ for $N$ odd. To be concrete, we focus on spin manifolds in this discussion. As an aside, the minimal TQFT plays a crucial role in the construction of non-invertible symmetry defects in 4d QFTs (see, e.g. Kaidi et al. (2022); Choi et al. (2023a)).
@@
where $\widehat{\underline{B}}_{2}\in H^{2}(Y^{4};\mathbb{Z})$. Equivalently, as was pointed out by Karasik (2023), the minimal TQFT $\mathcal{A}^{N,p}$ has a non-vanishing partition function only if the flux of the background connection through every 2-cycle of $X^{3}$ is a multiple of NN. [^22] A similar observation was made in Appendix A of Choi et al. (2023a).
@@
where the cohomology operation $\mathfrak{B}_{p}:H^{i}(-;\mathbb{Z}_{pm})\to H^{pi}(-;\mathbb{Z}_{p^{2}m})$ for any prime $p$ and any $m\in\mathbb{Z}^{+}$ is the Pontryagin $p$-th power Thomas (1956). The case of $p=2$ and $m=1$ corresponds to the standard Pontryagin square operation Whitehead (1949). One of its axioms is that
@@
where $\cup_{i}:C^{|x|}(-;\mathbb{Z}_{pm})\times C^{|y|}(-;\mathbb{Z}_{pm})\to C^{|x|+|y|-i}(-;\mathbb{Z}_{p^{2}m})$ denotes the cup-$i$ product Steenrod (1947), satisfying
@@ 4.2 Vanishing of the RR partition function in 2d
Consider a free Dirac fermion in $2$ dimensions. Like the $4d$ system, this has vector and axial $\mathrm{U}(1)$ symmetries with a mixed anomaly, leading to a vanishing condition similar to the one in the case of the ABJ anomaly discussed in the introduction. However, in this case, there is an additional structure that we can exploit, as discussed in Karch et al. (2019).
To define a fermion theory on a Riemann surface $\Sigma^{2}$, we need to specify a Spin structure. One can think of this as specifying periodic (R) or antiperiodic (NS) boundary conditions to each cycle on $\Sigma^{2}$. A classical result, derived e.g. in Hori et al. (2003), is that the partition function of a free fermion on the torus with RR Spin structure vanishes. When quantising the theory, one finds that there is a fermionic zero mode, which causes the path integral to vanish. On general $\Sigma^{2}$, the appearance of such zero modes, and therefore the vanishing, is controlled by the mod $2$ index, i.e. the number of zero modes modulo $2$. In $2d$, this index can be non-zero and is a topological invariant. A theorem due to Atiyah Atiyah (1971) states that
@@
where $\mathcal{I}(\rho_{\Sigma^{2}})$ is the mod $2$ index associated to the Spin structure $\rho_{\Sigma^{2}}$. To understand the RHS, recall that the first homology group with $\mathbb{Z}_{2}$ coefficients of the genus $g$ Riemann surface is given by
@@
We may define the intersection form for two cycles represented by closed curves $\gamma$, $\delta$ in $\Sigma^{2}$: [^24] We are making the additional technical assumption that $\gamma$ and $\delta$ only intersect transversely. Intuitively, we can always deform them slightly to satisfy this assumption without changing the homology classes they represent. A proof can be found in standard references for differential topology, such as Guillemin and Pollack (2010).
@@
for any $a,b\in H_{1}(\Sigma^{2};\mathbb{Z}_{2})$. It can be shown that there is a bijection between quadratic refinements of the intersection form and spin structures on $\Sigma^{2}$ Johnson (1980). Once we have such a quadratic refinement of the intersection form, we can define its associated $\mathbb{Z}_{2}$-valued Arf invariant,
@@
where $\gamma^{3}$ is the chirality matrixKarch et al. (2019); Delmastro et al. (2021). By general arguments Witten and Yonekura (2020), the anomaly theory of this system is given by the $\eta$ invariant of a real Dirac operator in $3d$ coupled to this $\mathbb{Z}_{2}$-background. Potential anomalies of this symmetry are classified by $\Omega^{\text{Spin}}_{3}(B\mathbb{Z}_{2})=\mathbb{Z}_{8}$. This group is generated by $\mathbb{RP}^{3}$ with non-trivial $\mathbb{Z}_{2}$ bundle. As pointed out in Tachikawa and Yonekura (2019), the relevant $\eta$-invariant evaluates to $\pm 1/8$ on this generator, so our $(-1)^{F_{L}}$-symmetry is indeed anomalous. To see the connection between this $\eta$-invariant and the Arf invariant discussed above we need to take one additional mathematical detour.
A quadratic enhancement of the intersection form is a function $\tilde{q}:H_{1}(\Sigma^{2};\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{4}$ such that $\tilde{q}(a+b)=\tilde{q}(a)+\tilde{q}(b)+2a\cdot b$. A Pin- structure gives rise to such a quadratic enhancement. Since every closed $2$-manifold admits a Pin- structure Kirby and Taylor (1991) but only orientable surfaces admit a Spin structure, this provides a generalisation of the quadratic refinement above. On spin surfaces, the quadratic enhancement $\tilde{q}$ is related to the quadratic refinement $q$ as
@@
We omit further details of this construction and refer the interested reader to Kirby and Taylor (1991).
@@
This theory has been discussed in the physics literature, see e.g. Guo et al. (2020); Grigoletto and Putrov (2023). The Poincaré dual of $x$ can be represented by a closed surface embedded in $\mathcal{M}^{3}$. Since every closed surface is pin-, the ambient Spin structure $\rho$ induces a Pin- structure $s$ on PD(x)(x). Consider our generator of $\Omega^{\text{Spin}}_{3}(B\mathbb{Z}_{2})$. The Poincaré dual of the non-trivial element $x\in H^{1}(\mathbb{RP}^{3};\mathbb{Z}_{2})$ is an embedded $\mathbb{RP}^{2}$. Depending on the choice of Pin- structure, $\beta(\mathbb{RP}^{2},s)=\pm 1\in\mathbb{Z}_{8}$. Therefore, the partition function of the theory $\mathcal{A}$ agrees with the exponentiated $\eta$-invariant on the generator of $\Omega^{\text{Spin}}_{3}(B\mathbb{Z}_{2})$. Since both quantities are bordism invariants, we deduce that they agree on all closed spin $3$-manifolds with $\mathbb{Z}_{2}$ bundle. That is, $\mathcal{A}$ is a presentation of the anomaly theory of our Majorana fermion.